first clear of fractions then solve using substitution
x/2 - y/3= 5/6
x/5 - y/4= 29/10

Answer by Alan3354(69443)

first clear of fractions then solve using substitution
x/2 - y/3= 5/6
x/5 - y/4= 29/10
------------------
I wouldn't do it that way, but...
Multiply the 1st eqn by 6 & and the 2nd by 20
---
3x - 2y = 5
4x - 5y = 58
-------
Multiply the 1st eqn by 4 and the 2nd by 3
12x - 8y = 20 --> 12x = 8y+20
12x -15y = 174
-----
Sub for 12x in the 2nd eqn
(8y+20) - 15y = 174
-7y = 154
y = -22
etc